Two softshell turtle species, midland smooth (Apalona mutica mutica) and spiny (Apalone spinifera) both consume the same small fishes. If midland smooth softshells consume 11 small fishes and spiny softshells consume 23 small fishes, what is b, the competition coefficient for the impact of midland smooth softshells on spiny softshells?

Select one:

a. 2.09

b. 0.48

c. 3.48

d. 1.09

Two softshell turtle species, midland smooth (Apalona mutica mutica) and spiny (Apalone spinifera) both consume the same small fishes. Midland smooth softshells consume, on average, 14 small fishes per day while spiny softshells consume 22 small fishes on average per day. What is a, the competition coefficient that represents the effect of spiny softshells on midland smooth softshells?

Select one:

a. 0.57

b. 1.57

c. 0.64

d. 1.64
Answer #1

Competition coefficient is the competitive effect of one species on another. Alpha12 and Alpha21 are the competition coefficients

Alpha 12 (a) is the effect of species 2 on species 1

Alpha 21 (b) is the effect of species 1 on species 2

species 1 is the affected species and species 2 is the species causing the effect

1st ) Answer is option b - 0.48

species 1 - midland smooth softshells

species 2 - spiny softshells

b (alpha21) = effect of species 1 on species 2

= 11/23 = 0.478 = 0.48

0.48 is the impact of midland smooth softshells on spiny softshells.

Therefore the answer is option b. 0.48

2) The correct answer is option b - 1.57

species 1 - midland smooth softshells

species 2 - spiny softshells

a (alpha12) = effect of species2 on species 1

= 22/14 = 1.57

1.57 is the impact of spiny softshells on midland smooth softshells

therefore the answer is option b. 1.57